OpenStudy (anonymous):

3(2K-2)=-2(4k-11) solve

OpenStudy (anonymous):

6k-6=-8k+22 6k+8k=22+6 14k=28 k=2

OpenStudy (lilg132):

k = 2

OpenStudy (anonymous):

6k-6=-8k+22 6k+8k=22+6 (14k=28)x1/14 k=2 2 is the answer